{6x+4y=12
{2y=6-3x
graph and classify each system. Then find the solution from the graph

Answer by richwmiller(17219)

{6x+4y=12
{2y=6-3x
multiply the second by 2
{2y=6-3x
4y=12-6x
rearrange
6x+4y=12
equations one and two are multiples of each other and have the same solution set